OpenStudy (anonymous):

what is the square root of -27

OpenStudy (anonymous):

\[3i \sqrt{3}\]

OpenStudy (anonymous):

imaginary since u take root of negative

OpenStudy (anonymous):

Now, how do you figure that out?

OpenStudy (anonymous):

square root(-27) = 5.19615242 i

OpenStudy (anonymous):

OK, here's the steps. Whenever a square root is negative, you can get rid of the negative sign by putting i on the outside. So all you have left is the square root of 27, do it like this\[i\sqrt{9*3}\]\[i(\sqrt{9}*\sqrt{3})\]\[3i \sqrt{3}\]